PART 1: Assume $50,000 is paid for an asset in present time that yields annual cash...

PART 1:

Assume $50,000 is paid for an asset in present time that yields annual cash flow of $6,000 dollars per year in Years 1 through 10, as shown in the table below. At the end of Year 10 the salvage value of the asset is $40,000. Calculate the rate of return on this investment. (In other words, find the value of the interest rate i that makes the Net Present Value of the investment equal to zero - this is also called the Internal Rate of Return.)

A hint here is that you will need to solve this problem either by trial and error or using the IRR function in Excel.

Note: Please enter your answer as a number to two decimal places. For example, if your answer is 5%, you would enter 5.00, not 0.05.

C=-$50,000 I=$6000 I=$6000 I=$6000 ... I=$6000 L=$40,000
0 1 2 3 ... 10

C: Cost, I: Income, L: Salvage Value

PART 2:
Suppose that the nominal annual interest rate on an investment is 12%. Calculate the effective interest rate if compounding occurs CONTINUOUSLY AND MONTHLY
Answer #1

Part 1:

0-50000 6000 6000 6000 76000 6000 6000 IRR 10.79%) 01234567890

IRR is calculated using the IRR function, =IRR(values), values are from year 0 to year 10.

PART-2

Let's assume principal to be 100, then we can calculate its future value by using the FV function in excel,

=FV(12%/12,12,0,100)

=112.68

Effective interest is present value - future value = 112.68 -100 = 12.68%